# Forest-of-Thought: Scaling Test-Time Compute for Enhancing LLM Reasoning

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable abilities across various language tasks, but solving complex reasoning problems remains a challenge. While existing methods like Chain-of-Thought (CoT) and Tree-of-Thought (ToT) enhance reasoning by decomposing problems or structuring prompts, they typically perform a single pass of reasoning and may fail to revisit flawed paths, compromising accuracy. To address this, we propose a novel reasoning framework called Forest-of-Thought (FoT), which integrates multiple reasoning trees to leverage collective decision-making for solving complex logical problems. FoT utilizes sparse activation strategies to select the most relevant reasoning paths, improving both efficiency and accuracy. Additionally, we introduce a dynamic self-correction strategy that enables real-time error correction and learning from past mistakes, as well as consensus-guided decision making strategies to optimize correctness and computational resources. Experimental results demonstrate that the FoT framework, combined with these strategies, significantly enhances the reasoning capabilities of LLMs, enabling them to solve complex tasks with greater precision and efficiency.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing by demonstrating remarkable abilities across a wide range of language tasks. Leveraging vast datasets and complex architectures, LLMs such as ChatGPT (Kojima et al., 2022; Achiam et al., 2023) and LLaMA (Touvron et al., 2023) can generate coherent essays, answer complex questions, and even engage in multi-turn dialogues with human-like fluency. These models excel at tasks requiring not only linguistic understanding but also basic reasoning, such as translating text between languages, summarizing lengthy documents, and creating code based on plain language instructions. The versatility and adaptability of LLMs have made them invaluable tools in both industry and research, simultaneously providing new avenues for addressing general-purpose problems.
@@
How to enable LLM to successfully solve hard reasoning problems is still challengable. A seires of works have been proposed to introduce more inference at test time based on a well-trained LLM (Wei et al., 2022; Yao et al., 2024; Snell et al., 2024; OpenAI, 2024). Chain-of-Thought (CoT) (Wei et al., 2022) provides a few chain of thought demonstrations in prompting as exemplars to emerge reasoning abilities of LLMs. Tree-of-Thought (ToT) (Yao et al., 2024) allowing language models to explore multiple reasoning paths and self-evaluate to make more globally informed decisions. Graph-of-Thought (GoT) (Besta et al., 2024) advances LLM prompting by structuring information as a graph of interconnected “thoughts”, enabling synergistic reasoning and feedback loops.
@@
These method perform reasoning by richer prompt or decomposing complex problem into several easier subproblems. They only perform a single complete reasoning pass on the problem, which cannot ensure that the problem is solved or guarantee correctness. For example, in a complex mathematical word problem, a Tree-of-Thought approach might decompose the problem into smaller steps, such as isolating terms or simplifying expressions. However, while breaking down the question, it may still overlook critical details or make errors in intermediate steps, leading to an incorrect final answer. Once it completes a single reasoning path, it typically does not revisit other possible approaches if the initial path is flawed. This lack of re-evaluation can result in a solution that fails to address the full complexity of the problem, as alternative paths are often prematurely abandoned and left unexplored, thereby compromising accuracy. Instead, humans tend to repeatedly think and verify from different perspectives when dealing with complex problems to truly solve the problem and provide answers with higher accuracy.
@@
In this paper, we propose a new reasoning framework named Forest-of-Thought (FoT) to scale up test-time compute for enhancing the reasoning abilities of LLMs, as shown in Figure 1. FoT integrates multiple reasoning trees to leverage the advantages of collective decision-making in handling complex logical reasoning tasks. By utilizing sparse activation strategies, we select the most relevant reasoning paths for each tree, improving both the efficiency and accuracy of the model. To further enhance the reasoning process, we propose a dynamic self-correction strategy, which allows the model to automatically identify and correct errors during the reasoning process, leveraging both real-time correction and historical learning. Additionally, we incorporate consensus-guided decision making strategies to optimize the correctness computational resources, ensuring that the model only continues the reasoning process when necessary. Our experiments demonstrate that the proposed FoT framework, combined with these strategies, significantly improves the reasoning performance of LLMs, enabling them to solve complex tasks with greater precision and efficiency.
@@ 2.1 XoT reasoning
Starting from Chain-of-Thought (CoT), XOT reasoning (e.g., ToT and GoT) began to be an important technique to enhance the reasoning ability of LLMs, leading to the development of a series of XoT reasoning algorithms.
@@
Chain-of-Thought (CoT) (Wei et al., 2022) decomposes a problem into a series of intermediate steps, each of which provides a portion of the information needed for the final answer. This approach mimics human problem-solving strategies, involving step-by-step reasoning to reach a conclusion. However, while the CoT method performs well in many tasks, it has limitations when dealing with complex mathematical and logical problems. These intricate problems often require multidimensional, non-linear thinking, rather than just sequential reasoning. Despite numerous subsequent studies that have improved the CoT, including Zero-Shot-CoT (Kojima et al., 2023), Self-Consistency with CoT (CoT-SC) (Wang et al., 2023), Auto-CoT (Zhang et al., 2022), VerifyCoT (Zhao et al., 2023), CoF-CoT (Nguyen et al., 2023), further exploration and optimization are still necessary to address highly complex tasks.
@@
Least-to-Most Prompting (LtM) (Zhou et al., 2023) leads the model through a step-by-step process, progressively assisting it in constructing a solution, in contrast to methods like CoT that attempt to solve complex problems directly. This approach effectively avoids the reasoning errors that can arise when attempting to solve complex problems all at once. By decomposing complex problems into a series of simpler tasks, Program of Thought (PoT) (Chen et al., 2023), Chain of Code (CoC) (Li et al., 2024) and Buffer of Thought (BoT) (Yang et al., 2024) transform this process into a set of programmatic steps, using variable names to convey semantic information. On the other hand, Algorithm of Thought (AoT) (Sel et al., 2024) method aims to integrate these steps into a single prompt, enabling LLMs to learn how to break down problems, generate solutions, assess their feasibility, and determine the next step in the search process. This approach reduces token consumption and improves efficiency.
@@
Tree-of-Thought (ToT) (Yao et al., 2024) constructs a tree structure to explore various possible choices and their outcomes, where each node represents a decision point and the edges represent transitions from one state to another. Typically, a depth-first search (DFS) approach is used to gradually explore each branch. Tree Prompting (Morris et al., 2023) establishes a decision-tree-based prompting system, chaining multiple language model calls together to collaboratively complete a specific task. However, for complex problems, the depth of the tree may become very large, leading to an exponential increase in the search space and an increased computational burden. Graph-of-Thought (GoT) (Besta et al., 2024) builds upon the ToT framework by introducing an aggregation process. GoT models the reasoning process of LLMs as a graph structure, allowing information units to form arbitrary dependencies, not limited to linear or tree-based arrangements. Through the aggregation process, GoT can consolidate information from multiple paths, supporting complex dynamic path selection and backtracking. Skeleton-of-Thought (SoT) (Ning et al., 2023) reduces generation latency in large language models (LLMs) by first generating a skeleton answer outline, then completing content in parallel, achieving significant speed-ups and potential quality improvements across various question types.
@@ 2.2 Monte Carlo Tree Search
Monte Carlo Tree Search (MCTS) (Chaslot et al., 2008), a probability-based search algorithm, has achieved significant progress across various domains since its introduction in computer Go in 2006. It evaluates nodes through randomized simulations (rollouts) and incrementally builds a local game tree to identify optimal or near-optimal solutions within limited time. To enhance MCTS performance, researchers have proposed various improvements.  (Browne et al., 2012) surveyed extensions like sparse activation, dynamic pruning, parallelization, and distributed computation, broadening MCTS applications.  (Srinivas et al., 2012) introduced UCB1-Tuned, an improved exploration-exploitation strategy suited for high-dimensional spaces. Recently, integrating MCTS with large language models has advanced its use in complex reasoning tasks. Methods like MCTSr (Zhang et al., 2024) combine MCTS with LLMs to guide decision-making, blending probabilistic search with linguistic reasoning. This hybrid approach enables effective multi-step logic and diverse solution paths.
@@ 3 Method
By introducing multiple reasoning trees (e.g., ToT (Yao et al., 2024) or MCTSr (Zhang et al., 2024)) for independent decision-making and employing sparse activation strategies to filter the results of key trees, we can construct an integrated framework known as the “forest of thought” to enhance the reasoning capability of LLMs, as shown in Figure 1(d) and Algorithm 1. This strategy leverages collective wisdom to compensate for individual deficiencies, thereby enhancing the model’s ability to think comprehensively from multiple directions. Our experimental validation demonstrates that the practice of integrating the results of multiple reasoning trees through sparse activation strategies indeed enhances the reasoning capabilities of larger models to a certain extent. This discovery not only enriches our understanding of model integration methods but also offers novel insights and approaches for improving the mathematical reasoning abilities of large language models (LLMs).
@@ 3.1 The FoT Framework
Suppose we have $n$ reasoning trees $T_{1},T_{2},\cdots,T_{n}$, each of which approaches the input problem from a different perspective. The root node of each tree represents the initial state or problem input, and each node represents an intermediate result or step in the reasoning process. Let the input be $x$. Each reasoning tree will start from the input and produce a result through different reasoning steps:
@@
where $\varepsilon(x)$ is the function that enhances $x$ before it is used as input (details in the following paragraph) and $T_{i}(\cdot)$ represents the reasoning process of the $i$th tree. FoT will consider the results of these trees in a sparse activation manner and produce high-quality response via majority voting. A dynamic self-correction strategy is proposed for enhancing the accuracy (Sec. 3.2), and early stopping strategies are introduced for improving inference efficiency (Sec. 3.3).
Sparse Activation.
@@
In the reasoning process of a forest of reasoning trees, only the most relevant reasoning trees (or certain nodes within them) are selected for computation, rather than performing a complete calculation on all trees. This approach enhances efficiency while improving model accuracy by selecting the most relevant reasoning paths. Through sparse activation, we filter the activations of each reasoning tree, ensuring that only the paths of certain reasoning trees are “activated” for inference.
@@
For each reasoning tree $T_{i}$, each layer considers the top-scoring nodes for further reasoning. Nodes with the highest scores are selected to split into child nodes at the next layer. If the nodes at a certain level of the tree cannot produce valid outputs, the tree’s splitting process will terminate early, and the activation indicator value will be set to 0. Otherwise, the tree will continue splitting until the specified depth, and the activation indicator value will be set to 1. Formally, the activation indicator $\varphi_{i}$ for the reasoning tree $T_{i}$ can be represented as
@@
$\varphi_{i}=\begin{cases}1,&\text{if tree }T_{i}\text{ is activated}\\ 0,&\text{if tree }T_{i}\text{ is not activated}\end{cases}$ (2)
For each tree $T_{i}$, node-level activation is determined at each layer. Let the activation score of a node $k$ at layer $l$ of tree $T_{i}$ be $s_{i,l,k}$. A node is activated if its score is among the top $m$ scores at that layer:
$\mathcal{A}_{i,l}=\{k\mid s_{i,l,k}\text{ is in the top }m\text{ at layer }l\}$ (3)
where $\mathcal{A}_{i,l}$ is the set of activated nodes at layer $l$ for tree $T_{i}$.
The output of the forest, incorporating both tree-level and node-level activations, is defined as:
$y=\sum_{i=1}^{n}\varphi_{i}\cdot\left(\sum_{l=1}^{L_{i}}\sum_{k\in\mathcal{A}_{i,l}}f_{i,l,k}(x)\right),$ (4)
where the total number of trees in the forest is indicated by $n$. The total number of layers in the tree $T_{i}$ is indicated by $L_{i}$. In tree $T_{i}$, $f_{i,l,k}(x)$ denotes the reasoning function of node $k$ at layer $l$. The set of activated nodes at layer $l$ in tree $T_{i}$ is indicated by $\mathcal{A}_{i,l}$.
Input Data Augmentation.
@@
When confronted with complex problems, our cognitive processes often transition from rapid and intuitive “fast thinking” to more in-depth and systematic “slow thinking”. This shift serves not only to address immediate challenges but also to leverage the relevant prior knowledge stored in our brains, enabling a more comprehensive analysis and resolution of problems. The slow thinking process integrates and evaluates this prior knowledge, allowing us to examine the problem from multiple perspectives and arrive at more reasonable solutions.
Inspired by this efficient problem-solving approach in humans, we have adopted a similar methodology to enhance the problem-solving capabilities of large language models. Unlike traditional “few-shot learning”, which merely presents a few examples for the model to imitate, this approach emphasizes extracting the most relevant portions from the model’s extensive knowledge base to assist its understanding and problem-solving. Specifically, when presented with a question or task $x$, the model initially attempts to find an answer through responding quickly. If this process fails to yield a satisfactory solution, the model initiates a deeper cognitive mechanism, filtering out highly relevant background information and prior experiences from a knowledge base $\mathcal{B}$ with several pre-collected samples:
$i_{max}=\arg\max_{i}\left(Sim\left(\mathrm{TF}(x),\mathrm{TF}(\mathcal{B}_{i})\right)\right)$ (5)
$x=x\oplus\mathcal{B}_{i_{max}}$ (6)
where $TF(\cdot)$ represents the TF-IDF transformation of the text which is fast yet effective, $Sim(\cdot)$ computes the cosine similarity between two vectors, and $\oplus$ means the concatenation of two text strings. This process mirrors human slow thinking, where the integration and analysis of pertinent information enable the model to construct a more complex cognitive framework, thereby enhancing its problem-solving capabilities and accuracy.
@@
4:     if activator indicator $\varphi_{i}==1$ then
5:        Dynamic self-correction:$s^{\prime}_{i}\leftarrow self\text{-}correct(s_{i},x)$ (Sec. 3.2);

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable abilities across various language tasks, but solving complex reasoning problems remains a significant challenge. While existing methods, such as Chain-of-Thought (CoT) and Tree-of-Thought (ToT), enhance reasoning by decomposing problems or structuring prompts, they typically perform a single pass of reasoning and may fail to revisit flawed paths, compromising accuracy. To address this limitation, we propose a novel reasoning framework called Forest-of-Thought (FoT), which integrates multiple reasoning trees to leverage collective decision-making for solving complex logical problems. FoT employs sparse activation strategies to select the most relevant reasoning paths, improving both efficiency and accuracy. Additionally, we introduce a dynamic self-correction strategy that enables real-time error correction, along with consensus-guided decision-making strategies to optimize both correctness and computational resources. Experimental results demonstrate that the FoT framework, combined with these strategies, significantly enhances the reasoning capabilities of LLMs, enabling them to solve complex tasks with greater precision and efficiency.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing by demonstrating remarkable capabilities across a wide range of language tasks. By leveraging vast datasets and complex architectures, LLMs such as ChatGPT (Kojima et al., 2022; Achiam et al., 2023) and LLaMA (Touvron et al., 2023) can generate coherent essays, answer complex questions, and engage in multi-turn dialogues with human-like fluency. These models excel at tasks requiring not only linguistic understanding but also basic reasoning, such as translating text, summarizing lengthy documents, and generating code from plain language instructions. The versatility and adaptability of LLMs have made them invaluable tools in both industry and research, opening up new avenues for solving general-purpose problems.
@@
Enabling LLMs to successfully solve complex reasoning problems remains a challenge. A series of works have been proposed to introduce more inference during testing based on a well-trained LLM (Wei et al., 2022; Yao et al., 2024; Snell et al., 2024; OpenAI, 2024). Chain-of-Thought (CoT) (Wei et al., 2022) provides a few chain-of-thought demonstrations in prompting as exemplars to enhance the reasoning abilities of LLMs. Tree-of-Thought (ToT) (Yao et al., 2024) allows language models to explore multiple reasoning paths and self-evaluate to make more globally informed decisions. Graph-of-Thought (GoT) (Besta et al., 2024b) advances LLM prompting by structuring information as a graph of interconnected “thoughts”, enabling synergistic reasoning and feedback loops.
@@
These methods perform reasoning by using richer prompts or by decomposing a complex problem into several simpler sub-problems. However, they typically perform only a single, complete reasoning pass on the problem, which does not guarantee that the problem will be solved correctly. For example, in a complex mathematical word problem, a Tree-of-Thought approach might decompose the problem into smaller steps, such as isolating terms or simplifying expressions. However, while breaking down the problem, it may still overlook critical details or make errors in intermediate steps, leading to an incorrect final answer. Once it completes a single reasoning path, it typically does not revisit other possible approaches if the initial path is flawed. This lack of re-evaluation can result in a solution that fails to address the full complexity of the problem, as alternative paths are often prematurely abandoned and left unexplored, thereby compromising accuracy. In contrast, humans tend to repeatedly reflect and verify from different perspectives when dealing with complex problems, which allows them to truly solve the problem and provide answers with higher accuracy.
@@
In this paper, we propose a new reasoning framework called Forest-of-Thought (FoT) to scale up test-time computation and enhance the reasoning abilities of LLMs, as shown in Figure 1. FoT integrates multiple reasoning trees to leverage the advantages of collective decision-making for handling complex logical reasoning tasks. By utilizing sparse activation strategies, we select the most relevant reasoning paths for each tree, thereby improving both the efficiency and accuracy of the model. To further enhance the reasoning process, we introduce a dynamic self-correction strategy, which enables the model to automatically identify and correct errors during reasoning, leveraging both real-time corrections and historical learning. Additionally, we incorporate consensus-guided decision-making strategies to optimize both correctness and computational resource usage, ensuring that the model only continues the reasoning process when necessary. Our experiments demonstrate that the proposed FoT framework, combined with these strategies, significantly improves the reasoning performance of LLMs, enabling them to solve complex tasks with greater precision and efficiency.
@@ 2.1 XoT reasoning
Starting with Chain-of-Thought (CoT), XoT reasoning techniques (e.g., ToT and GoT) emerged as important methods for enhancing the reasoning abilities of LLMs, leading to the development of a series of XoT reasoning algorithms.
@@
Chain-of-Thought (CoT) (Wei et al., 2022) decomposes a problem into a series of intermediate steps, each providing a portion of the information necessary for the final answer. This approach mimics human problem-solving strategies, involving step-by-step reasoning to reach a conclusion. However, while CoT performs well in many tasks, it has limitations when applied to complex mathematical and logical problems. These types of problems often require multidimensional, non-linear thinking rather than just sequential reasoning. Despite numerous subsequent studies aimed at improving CoT, such as Zero-Shot-CoT (Kojima et al., 2023), Self-Consistency with CoT (CoT-SC) (Wang et al., 2023), Auto-CoT (Zhang et al., 2022), VerifyCoT (Zhao et al., 2023), and CoF-CoT (Nguyen et al., 2023), further exploration and optimization are still needed to tackle highly complex tasks.
@@
Least-to-Most Prompting (LtM) (Zhou et al., 2023) guides the model through a step-by-step process, progressively assisting it in constructing a solution, in contrast to methods like CoT, which attempt to solve complex problems directly. This approach effectively mitigates the reasoning errors that often occur when attempting to solve complex problems all at once. By decomposing complex problems into simpler tasks, Program of Thought (PoT) (Chen et al., 2023), Chain of Code (CoC) (Li et al., 2024), and Buffer of Thought (BoT) (Yang et al., 2024) transform the process into a set of programmatic steps, using variable names to convey semantic information. In contrast, the Algorithm of Thought (AoT) (Sel et al., 2024) method seeks to integrate these steps into a single prompt, enabling LLMs to learn how to break down problems, generate solutions, assess their feasibility, and determine the next step in the search process. This approach reduces token consumption and improves efficiency.
@@
Tree-of-Thought (ToT) (Yao et al., 2024) constructs a tree structure to explore various possible choices and their outcomes, where each node represents a decision point, and the edges represent transitions between states. Typically, a depth-first search (DFS) approach is used to explore each branch incrementally. Tree Prompting (Morris et al., 2023) establishes a decision-tree-based prompting system, chaining multiple language model calls together to collaboratively complete a specific task. However, for complex problems, the tree’s depth may become very large, leading to an exponential increase in the search space and a higher computational burden. Graph-of-Thought (GoT) (Besta et al., 2024b) extends the ToT framework by introducing an aggregation process. GoT models the reasoning process of LLMs as a graph structure, allowing information units to form arbitrary dependencies, not limited to linear or tree-based arrangements. Through aggregation, GoT consolidates information from multiple paths, enabling complex dynamic path selection and backtracking. Skeleton-of-Thought (SoT) (Ning et al., 2023) reduces generation latency in LLMs by first generating a skeleton outline of the answer, then completing the content in parallel, achieving significant speedups and potential quality improvements across various types of questions.
@@ 2.2 Monte Carlo Tree Search
Monte Carlo Tree Search (MCTS) (Chaslot et al., 2008), a probability-based search algorithm, has made significant progress across various domains since its introduction in computer Go in 2006. It evaluates nodes through randomized simulations (rollouts) and incrementally builds a local game tree to identify optimal or near-optimal solutions within a limited time. To enhance MCTS performance, researchers have proposed various improvements. Browne et al. (2012) surveyed extensions such as sparse activation, dynamic pruning, parallelization, and distributed computation, which have broadened MCTS applications. Srinivas et al. (2012) introduced UCB1-Tuned, an improved exploration-exploitation strategy suited for high-dimensional spaces. Recently, integrating MCTS with large language models has advanced its use in complex reasoning tasks. Everything-of-Thought (Ding et al., 2024) employs a combination of reinforcement learning and MCTS for reasoning. Methods like MCTSr (Zhang et al., 2024) combine MCTS with LLMs to guide decision-making, blending probabilistic search with linguistic reasoning. This hybrid approach enables effective multi-step logic and diverse solution paths.
@@ 3 Method
By introducing multiple reasoning trees (e.g., ToT (Yao et al., 2024) or MCTSr (Zhang et al., 2024)) for independent decision-making and employing sparse activation strategies to filter the results from key trees, we can construct a Forest of Thought to enhance the reasoning capabilities of LLMs, as shown in Figure 1(d) and Algorithm 1. This strategy leverages collective intelligence to compensate for individual shortcomings, thereby improving the model’s ability to reason from multiple perspectives. Our experimental validation demonstrates that integrating the results from multiple reasoning trees through sparse activation strategies indeed enhances the reasoning capabilities of larger models to a significant extent. This finding not only deepens our understanding of model integration techniques but also provides new insights for improving the mathematical reasoning abilities of LLMs.
@@ 3.1 The FoT Framework
Suppose we have $n$ reasoning trees $T_{1},T_{2},\cdots,T_{n}$, each of which approaches the input problem from a different perspective. The root node of each tree represents the initial state or problem input, and each subsequent node represents an intermediate result or step in the reasoning process. Let the input be $x$. Each reasoning tree will start from the input and produce a result through different reasoning steps:
@@
where $\varepsilon(x)$ is a function that enhances $x$ before it is used as input (details in the following paragraph) and $T_{i}(\cdot)$ represents the reasoning process of the $i$th tree. FoT considers the results of these trees in a sparse activation manner and produces a high-quality response through a decision-making strategy (Sec. 3.3). Additionally, a dynamic self-correction strategy is proposed to enhance the accuracy (Sec. 3.2).
@@
Sparse Activation. In the reasoning process of forest of thoughts, we improve computational efficiency and accuracy by selecting only the most relevant reasoning trees for computation, rather than performing full calculations on all trees. Through sparse activation, we filter which reasoning trees are activated, ensuring that only the paths of selected trees are used for inference.
@@
For each reasoning tree $T_{i}$, at each layer, we select the top-scoring nodes for further reasoning. The highest-scoring nodes are chosen and split into child nodes in the next layer. If the nodes at a given level cannot produce valid outputs, the tree’s splitting process terminates early, and the activation indicator $\varphi_{i}$ is set to 0. Otherwise, the tree continues splitting until the specified depth, and the activation indicator $\varphi_{i}$ is set to 1. Formally, the activation indicator $\varphi_{i}$ for reasoning tree $T_{i}$ is defined as:
@@
$\varphi_{i}=\begin{cases}1,&\text{if }\forall l,\,F(s_{l})=\text{valid output},\\ 0,&\text{otherwise}.\end{cases}$ (2)
where $\varphi_{i}$ is the activation indicator of the $i$-th reasoning tree. $s_{l}$ is the solution of the $l$-th layer node in the reasoning tree, and $F(s_{l})$ is the function used to extract valid output from the node $s_{l}$. A tree is considered active if, for every layer $l$, the function $F$ returns a valid output. If any layer fails to produce a valid output, the tree’s reasoning process is stopped, and the activation indicator $\varphi_{i}$ is set to 0, indicating that the tree is inactive. Ultimately, only active trees participate in the decision-making process of the FoT’s result.
@@
Input Data Augmentation. When faced with complex problems, our cognitive process typically shifts from rapid, intuitive “fast thinking” to deeper, more systematic “slow thinking” (Evans, 1984; Kahneman, 2011). This transition not only helps us address immediate challenges but also draws on the relevant prior knowledge stored in our brains, enabling us to analyze and solve problems more comprehensively. Slow thinking integrates and evaluates this prior knowledge, allowing us to view problems from multiple perspectives and arrive at more reasonable solutions. Inspired by prior knowledge, we have collected and constructed a knowledge base $\mathcal{B}$ from publicly available datasets to support the model’s reasoning process. Additionally, by enhancing the diversity of question-answer generation within each tree to further improve the model’s performance.
$i_{max}$ $=\arg\max_{i}\left(Sim\left(x,\mathcal{B}\right)\right)$ (3)
$x$ $=\mathcal{B}_{i_{max}}\oplus x$ (4)
where $Sim(\cdot)$ represents retrieving the most relevant questions from $\mathcal{B}$. $\oplus$ means the concatenation of two text strings.
The sparse activation strategy improves both efficiency and accuracy by focusing on the most relevant reasoning paths and reducing unnecessary computations. To further optimize reasoning, we apply efficient search with early termination. When a clear target is identified (e.g., Game of 24, Arithmetic Puzzles and Maze Solving), the search halts immediately upon finding a solution. This avoids redundant computations: once a branch matches the ground truth (GT), further exploration of irrelevant paths is stopped, saving resources and speeding up the process. Early termination thus enhances overall efficiency by preventing unnecessary work.
@@
4:     dynamic self-correction:$s^{\prime}_{i}\leftarrow self\text{-}correct(s_{i},x)$ (Sec. 3.2);
5:     if activator indicator $\varphi_{i}==1$ then